OpenStudy (anonymous):

Can anyone help me set up an optimization problem? A box with a square base and open top must have a volume of 32,000 cm . Find the dimensions of the box that minimize the amount of material used.

OpenStudy (anonymous):

So I know that b/c we have a square base, we're looking at x^2*height=32,000cm^3

OpenStudy (anonymous):

I am stuck though, I can not set up the f(x) equation. My brain just isn't seeing it.

hartnn (hartnn):

amount of material used. --- > total surface area = ... ?

OpenStudy (anonymous):

I don't see what you're getting at?

OpenStudy (anonymous):

so height * width * length = volume

hartnn (hartnn):

thats correct, but you have to minimize amount of material used. --- > minimize total surface area

OpenStudy (anonymous):

amount of material might be just h * w * 4, + w * l?

OpenStudy (anonymous):

sure, I have to minimize my total surface area, yes

hartnn (hartnn):

thats correct, with w=l

OpenStudy (anonymous):

yes!, how do I set up my equation?

OpenStudy (anonymous):

let x = width and length

OpenStudy (anonymous):

32000=x^2 * l

OpenStudy (anonymous):

I am NOT seeing this.

hartnn (hartnn):

f(x) = 4 h x + x^2 [total surface area] volume = 32000 = x^2 h got this ? h= height.

OpenStudy (anonymous):

Height, that's right... still processing in my brain, though, sorry.

OpenStudy (anonymous):

okay, I get the f(x), but am I tossing out my volume?

hartnn (hartnn):